Many investors and financial analysts believe the Dow Jones Industrial Average (DJIA) gives a good barometer of the overall stock market. On January 31, 2006, 9 of the 30 stocks making up the DJIA increased in price (The Wall Street Journal, February 1, 2006). On the basis of this fact, a financial analyst claims we can assume that 30% of the stocks traded on the New York Stock Exchange (NYSE) went up the same day.
A sample of 57 stocks traded on the NYSE that day showed that 23 went up.
You are conducting a study to see if the proportion of stocks that went up is is significantly more than 0.3. You use a significance level of 0.01
What is the test statistic for this sample? (Report answer accurate to three decimal places.)
test statistic
What is the p-value for this sample? (Report answer accurate to four decimal places.)
p-value
.

Answer by Boreal(15235)

the test statistic is (p hat-0.3)/sqrt (0.3*0.7/57) and is a z; p hat is 23/57=0.4035
if z>2.326, then reject Ho
test stat is z=(0.4035-0.30)/sqrt(0.21/57)
=0.1035/0.0607
=1.705
p-value is z>1.705 with p-value of 0.0441, which is greater than 0.01